# Regge trajectories and bridges between them in integrable AdS/CFT

### v2

@@ Appendix C Proof of some QSC constraints in special sector
$F^{[\pm n]}(u)\equiv F\left(u\pm\frac{in}{2}\right),$ (188)
@@
$=\mathbb{R}_{a}^{c}{\bf P}_{cr}^{[+1]}(u)\kappa^{gq}Q_{q|p}(u)\,\mathbb{K}_{j}^{p}\Theta_{i}^{j\,[+2]}(u).$
Now, using expression (189), we can compare (192) and (LABEL:second). We get
@@
Proof $\Theta$ asymptotics.
@@
$(\mathbb{R}^{-1})^{b}_{a}e^{\mathrm{i}\pi(\mathcal{N}_{b}+\hat{\mathcal{N}}_{i})}|u|^{\mathcal{N}_{b}+\hat{\mathcal{N}}_{i}}B_{b|i}^{\downarrow}=|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{j}}\Theta_{i}^{j}(u)B_{a|i}^{\downarrow},$ (198)
$(\mathbb{R}^{-1})^{a}_{a}e^{\mathrm{i}\pi(\mathcal{N}_{a}+\hat{\mathcal{N}}_{i})}|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}B_{a|i}^{\downarrow}=|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{j}}\Theta_{i}^{j}(u)B_{a|i}^{\downarrow},$ (199)
@@
as $u\rightarrow+\infty$ (in the last line and in LHS of the second line there is no sum of repeated indices). In the derivation we used the fact that $\mathbb{R}$ is diagonal together with (2.1.3), which imply that $\Theta$ is asymptotically diagonal. It seems that the LHS depends on index $a$ whereas the RHS does not, but one can note that this dependence disappears, indeed $(\mathbb{R}^{-1})_{a}^{a}\text{e}^{\mathrm{i}\pi\mathcal{N}_{a}}=1$ afor-$\forall a$. The $u\rightarrow+\infty$ asymptotics (72) immediately follows. To find the behaviour as $u\rightarrow-\infty$ one can use (70).
@@
From the definition of the gluing matrix we have $\Theta^{[-]}(u)=\mathcal{G}^{-1}(u)\cdot f(u)$ and substituting that in (71) one gets
@@
Now the goal is to rewrite $f^{[+2]}(u)\cdot\mathbb{K}\cdot f^{-1}(u)$ in a simpler way using the properties of the $\tau$-functions. Moreover we notice that $\mathbb{K}=\mathbb{K}^{T}=\mathbb{K}^{-1}$ and $[\kappa,\mathbb{K}]=0$. First of all it holds
$\tau^{j[+2]}=\kappa^{ji}\tau^{[+4]}_{i}=\kappa^{ji}\mathbb{K}_{i}^{l}\tau^{[+2]}_{l}=\mathbb{K}^{j}_{i}\kappa^{il}\tau^{[+2]}_{l}=\mathbb{K}^{j}_{i}\tau^{i}.$ (202)
@@
$\left(f^{[+2]}(u)\cdot\mathbb{K}\cdot f^{-1}(u)\right)_{i}^{j}$ $=(\delta_{i}^{l}-\mathbb{K}_{i}^{p}\tau_{p}\mathbb{K}^{l}_{r}\tau^{r})\mathbb{K}_{l}^{m}(\delta_{m}^{j}+\tau_{m}\tau^{j})$
$=(\mathbb{K}_{i}^{m}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{m})(\delta_{m}^{j}+\tau_{m}\tau^{j})$
$=\mathbb{K}_{i}^{j}+\mathbb{K}_{i}^{m}\tau_{m}\tau^{j}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{j}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{m}\tau_{m}\tau^{j}=\mathbb{K}_{i}^{j}\quad\Box.$ (203)
@@
By continuity from the case of local operator we can select the minus sign, see appendix below.
@@ Appendix D The case of local operators reviewed
We are considering the gluing matrix for local operators, hence $\mathcal{G}_{i}^{j}=\texttt{const}$. Due to LR symmetry, $\mathcal{G}=\mathbb{K}\cdot\mathcal{G}\cdot\mathbb{K}$. This fixes
$\mathcal{G}_{i}^{\,\,\,j}=\begin{pmatrix}g_{11}&0&0&g_{14}\\ 0&g_{22}&g_{23}&0\\ 0&g_{32}&g_{33}&0\\ g_{41}&0&0&g_{44}\\ \end{pmatrix}\,.$ (207)
Firstly, from definition (67), we can build the matrix $f$, and using the asymptotics of $\Theta(u)$ we deduce this should take the form
$\lim_{u\rightarrow+\infty}f_{i}^{j}(u)=\mathcal{G}_{i}^{\,\,\,k}\lim_{u\rightarrow+\infty}\Theta_{k}^{\,\,\,j}(u)=\begin{pmatrix}e^{\mathrm{i}\pi\hat{\mathcal{N}_{1}}}g_{11}&0&0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{4}}}g_{14}\\ 0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{2}}}g_{22}&e^{\mathrm{i}\pi\hat{\mathcal{N}_{3}}}g_{23}&0\\ 0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{2}}}g_{32}&e^{\mathrm{i}\pi\hat{\mathcal{N}_{3}}}g_{33}&0\\ e^{\mathrm{i}\pi\hat{\mathcal{N}_{1}}}g_{41}&0&0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{4}}}g_{44}\\ \end{pmatrix},$ (208)
and similarly at $u\rightarrow-\infty$ using (70). Now, we require the elementary property $f+f^{-1}=2\times\mathds{1}$, that is (69) in the symmetric sector,
$\mathcal{G}\cdot\lim_{u\rightarrow\pm\infty}\Theta\left(u\right)+\kappa\cdot\lim_{u\rightarrow\pm\infty}\Theta^{T}\left(u\right)\cdot\mathcal{G}^{T}\cdot\kappa^{-1}=2\times\mathds{1}.$
Finally, we impose the Q-system invariance $\mathcal{G}\cdot\kappa\cdot\mathcal{G}^{T}=\kappa$ together with the parity constraint $\mathcal{G}=-\kappa\cdot\mathcal{G}^{T}\cdot\kappa^{-1}$. Property $\text{det}(\mathcal{G})=1$ immediately follows.
  • Let $g_{32}\neq 0$, the above algorithm provides
    $\mathcal{G}_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)&0&0&\frac{\tan^{2}\left(\pi\Delta_{sl(2)}\right)}{g_{41}}\\ 0&\mathrm{i}\sec\left(\pi S_{sl(2)}\right)&\frac{\tan^{2}\left(\pi S_{sl(2)}\right)}{g_{32}}&0\\ 0&g_{32}&-\mathrm{i}\sec\left(\pi S_{sl(2)}\right)&0\\ g_{41}&0&0&\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)\\ \end{array}\right).$ (209)
  • Let $g_{32}=0$, the algorithm yields
    $\mathcal{G}_{i}^{\,\,\,j}=\left(\begin{array}[]{cccc}-\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)&0&0&\frac{\tan^{2}\left(\pi\Delta_{sl(2)}\right)}{g_{41}}\\ 0&\pm\mathrm{i}&g_{23}&0\\ 0&0&\mp\mathrm{i}&0\\ g_{41}&0&0&\mathrm{i}\sec\left(\pi\Delta_{sl(2)}\right)\\ \end{array}\right),\quad S_{sl(2)}\in\mathbb{Z}.$ (210)
Furthermore, we know that (see sections 2.1.3 and 3.1.1)
$\lim_{u\rightarrow\pm\infty}f_{i}^{\,\,\,j}(u)$ $=\left(\begin{array}[]{cccc}e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{4,\pm}+1&e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{3,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{2,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}^{2}\\ e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{4,\pm}&e^{\imath\mathcal{P}}t_{2,\pm}t_{3,\pm}+1&-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}^{2}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{2,\pm}\\ e^{\mathrm{i}\mathcal{P}}t_{3,\pm}t_{4,\pm}&e^{\mathrm{i}\mathcal{P}}t_{3,\pm}^{2}&1-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{3,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{3,\pm}\\ e^{\mathrm{i}\mathcal{P}}t_{4,\pm}^{2}&e^{\mathrm{i}\mathcal{P}}t_{3,\pm}t_{4,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{4,\pm}&1-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{4,\pm}\\ \end{array}\right)$ (215)
$\equiv\mathcal{G}_{i}^{\,\,\,k}\lim_{u\rightarrow\pm\infty}\Theta_{k}^{\,\,\,j}(u),$ (216)
@@
and we should assume that $\tau$ has constant asymptotics in order to have a gluing matrix independent of $u$.
The crucial step is now to use definition (215) to parametrize $\mathcal{G}$ by the charges and by the asymptotic values of $\tau_{i}(u)$. The two alternative expressions as $u\rightarrow\pm\infty$ should coincide, since $\mathcal{G}$ is a constant. This poses some additional constraints, and they can have only two possible outcomes: either $\Delta$ or $S$ needs to be quantized. In particular, by employing the asymptotic behavior of $\Theta(u)$ along with the prototype form (207) in (215), the conditions
@@
are immediately derived. Upon analyzing the solutions (209) and (210) within equation (215), it becomes evident that the case $t_{1,\pm}=t_{4,\pm}=0$ results in the quantization of $\Delta$. In order to avoid this restriction, one must choose $t_{2,\pm}=t_{3,\pm}=0$, which, through (215), forces to select the solution (210) with $g_{23}=0$ (and $S\in\mathbb{Z}$). Consequently, the gluing matrix (77), as reported in the main text, directly follows.
Appendix E Numerical results
@@
Table 1: Numerical results for the leading Regge trajectory interpolating the even spin operators in the $\Delta-S$ plane ($sl(2)$-like grading). Coupling is fixed to be $h=27/100$. In the numerical method, $N_{\text{trunc}}=22$. The error is estimated by truncating the last floating digit of the spin at the first iteration in which the error runs aground.
@@
$1.570786$ 0.40349193595508341712711865204630808300072004925(9)0.4034919359550834171271186520463080830007200492590.40349193595508341712711865204630808300072004925(9)0.40349193595508341712711865204630808300072004925 ( 9 )
$2.490786$ 1.15331291813913362657973321522920687329(6)1.1533129181391336265797332152292068732961.15331291813913362657973321522920687329(6)1.15331291813913362657973321522920687329 ( 6 )
$2.930786$ 1.53589583211995786945005245332082880571(9)1.5358958321199578694500524533208288057191.53589583211995786945005245332082880571(9)1.53589583211995786945005245332082880571 ( 9 )
$3.65$ 2.1804172586331162231220788403481454795192105310467872(2)2.180417258633116223122078840348145479519210531046787222.1804172586331162231220788403481454795192105310467872(2)2.1804172586331162231220788403481454795192105310467872 ( 2 )
$4.3$ 2.776703901183812605763361039625977959156428030(6)2.77670390118381260576336103962597795915642803062.776703901183812605763361039625977959156428030(6)2.776703901183812605763361039625977959156428030 ( 6 )
$4.95$ 3.3818980605041014213880577525181627895(1)3.381898060504101421388057752518162789513.3818980605041014213880577525181627895(1)3.3818980605041014213880577525181627895 ( 1 )
$5.7$ 4.088136216951466165858899839046551730378391365526927947(5)4.08813621695146616585889983904655173037839136552692794754.088136216951466165858899839046551730378391365526927947(5)4.088136216951466165858899839046551730378391365526927947 ( 5 )
Table 2: Numerical results for the Pomeron (“bottom” grading). The error is estimated by truncating the last floating digit of the spin at the first iteration in which the error runs aground.
$h$ $\mathcal{S}_{0}$ $N_{\text{trunc}}$
$0.005$ $1.00066520654474671981807717788039523366981$ $12$
80828574660352597718534181714(8)80828574660352597718534181714880828574660352597718534181714(8)80828574660352597718534181714 ( 8 )
$0.05$ $1.0565169798036575580020671070884020279449$ $12$
400322289775908178131(8)4003222897759081781318400322289775908178131(8)400322289775908178131 ( 8 )
$0.5$ 1.6736828615476622260025462071096(6)1.673682861547662226002546207109661.6736828615476622260025462071096(6)1.6736828615476622260025462071096 ( 6 ) $18$
$1$ 1.8274224988317420768251472449(4)1.827422498831742076825147244941.8274224988317420768251472449(4)1.8274224988317420768251472449 ( 4 ) $20$
$2$ 1.91194771316464271763438(4)1.9119477131646427176343841.91194771316464271763438(4)1.91194771316464271763438 ( 4 ) $24$
@@
Table 3: Numerical results for the minimum of the leading odd trajectory as a function of the coupling (“bottom” grading). The error is estimated by truncating the last floating digit of the spin at the first iteration in which the error runs aground.
$h$ $\mathcal{S}_{0}^{\text{odd}}$ $N_{\text{trunc}}$
$0.003$ 0.50018971502613513860323013063493040385095062082878530155(6)0.5001897150261351386032301306349304038509506208287853015560.50018971502613513860323013063493040385095062082878530155(6)0.50018971502613513860323013063493040385095062082878530155 ( 6 ) $12$
$0.01$ 0.5020944885214757093188076580147755273036194343998494360348(4)0.502094488521475709318807658014775527303619434399849436034840.5020944885214757093188076580147755273036194343998494360348(4)0.5020944885214757093188076580147755273036194343998494360348 ( 4 ) $12$
$0.1$ 0.63365211706208743377556493983082183433(5)0.6336521170620874337755649398308218343350.63365211706208743377556493983082183433(5)0.63365211706208743377556493983082183433 ( 5 ) $12$
$0.5$ 0.94753956996763001869486904196(2)0.9475395699676300186948690419620.94753956996763001869486904196(2)0.94753956996763001869486904196 ( 2 ) $18$
$1$ 0.978990578454073224755(4)0.97899057845407322475540.978990578454073224755(4)0.978990578454073224755 ( 4 ) $20$

### v3

@@ Appendix C Proof of some QSC constraints in special sector
$F^{[\pm n]}(u)\equiv F\left(u\pm\frac{\mathrm{i}n}{2}\right),$ (208)
@@
$=\mathbb{R}_{a}^{c}{\bf P}_{cr}^{[+1]}(u)\kappa^{gq}Q_{q|p}(u)\,\mathbb{K}_{j}^{p}\Theta_{i}^{j\,[+2]}(u).$ (215)
Now, using expression (209), we can compare (212) and (215). We get
@@
Proof $\Theta(u)$ asymptotics.
@@
$\sum_{b=1}^{4}(\mathbb{R}^{-1})^{b}_{a}e^{\mathrm{i}\pi(\mathcal{N}_{b}+\hat{\mathcal{N}}_{i})}|u|^{\mathcal{N}_{b}+\hat{\mathcal{N}}_{i}}B_{b|i}^{\downarrow}=\sum_{j=1}^{4}|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{j}}\Theta_{i}^{j}(u)B_{a|j}^{\downarrow},$ (219)
$(\mathbb{R}^{-1})^{a}_{a}e^{\mathrm{i}\pi(\mathcal{N}_{a}+\hat{\mathcal{N}}_{i})}|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{i}}B_{a|i}^{\downarrow}=\sum_{j=1}^{4}|u|^{\mathcal{N}_{a}+\hat{\mathcal{N}}_{j}}\Theta_{i}^{j}(u)B_{a|j}^{\downarrow},$ (220)
@@
as $u\rightarrow+\infty$ (the only sums over indices here are the ones explicitly written). In the derivation we used the fact that $\mathbb{R}$ is diagonal together with (2.1.3), which imply that $\Theta$ is asymptotically diagonal. It seems that the LHS depends on index $a$ whereas the RHS does not, but one can note that this dependence disappears, indeed $(\mathbb{R}^{-1})_{a}^{a}\text{e}^{\mathrm{i}\pi\mathcal{N}_{a}}=1$ afor-$\forall a$. The $u\rightarrow+\infty$ asymptotics (83) immediately follows. To find the behaviour as $u\rightarrow-\infty$ one can use (81).
@@
From the definition of the gluing matrix we have $\Theta^{[-1]}(u)=\mathcal{G}^{-1}(u)\cdot f(u)$ and substituting that in (82) one gets
@@
Now the goal is to rewrite $f^{[+2]}(u)\cdot\mathbb{K}\cdot f^{-1}(u)$ in a simpler way using the properties of the $\tau$-functions. Moreover, we notice that $\mathbb{K}=\mathbb{K}^{T}=\mathbb{K}^{-1}$ and $[\kappa,\mathbb{K}]=0$. First of all, using (29) it holds
$\tau^{j[+2]}=e^{-i\mathcal{P}}\kappa^{ji}\tau^{[+4]}_{i}=e^{-i\mathcal{P}}\kappa^{ji}\mathbb{K}_{i}^{l}\tau^{[+2]}_{l}=e^{-i\mathcal{P}}\mathbb{K}^{j}_{i}\kappa^{il}\tau^{[+2]}_{l}=\mathbb{K}^{j}_{i}\tau^{i}.$ (223)
@@
$\left(f^{[+2]}(u)\cdot\mathbb{K}\cdot f^{-1}(u)\right)_{i}^{j}$ $=(\delta_{i}^{l}-\mathbb{K}_{i}^{p}\tau_{p}\mathbb{K}^{l}_{r}\tau^{r})\mathbb{K}_{l}^{m}(\delta_{m}^{j}+\tau_{m}\tau^{j})$ (224)
$=(\mathbb{K}_{i}^{m}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{m})(\delta_{m}^{j}+\tau_{m}\tau^{j})$ (225)
$=\mathbb{K}_{i}^{j}+\mathbb{K}_{i}^{m}\tau_{m}\tau^{j}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{j}-\mathbb{K}_{i}^{p}\tau_{p}\tau^{m}\tau_{m}\tau^{j}=\mathbb{K}_{i}^{j}\quad\Box,$ (226)
which together with (222) proves the result.
@@
We observe here that the choice of the minus sign in (229) is fully consistent with the other constraints involved in the algorithm 3.2.1 presented in the main text. In contrast, the plus sign is incompatible with these constraints and prevents the algorithm from admitting a solution.
@@ Appendix D The case of local operators reviewed
We are considering the gluing matrix for local operators, hence $\mathcal{G}_{i}^{j}=\texttt{const}$. Due to LR symmetry and parity, $\mathcal{G}=\mathbb{K}\cdot\mathcal{G}\cdot\mathbb{K}$ and $\mathcal{G}=-\kappa\cdot\mathcal{G}^{T}\cdot\kappa^{-1}$. This fixes
$\mathcal{G}_{i}^{\,\,\,j}=\begin{pmatrix}g_{11}&0&0&g_{14}\\ 0&g_{22}&g_{23}&0\\ 0&g_{32}&-g_{22}&0\\ g_{41}&0&0&-g_{11}\\ \end{pmatrix}\,.$ (230)
Firstly, from definition (78), we can build the matrix $f(u)$, and using the asymptotics of $\Theta(u)$ (83), we deduce this should take the form
$\lim_{u\rightarrow+\infty}f_{i}^{\,\,\,j}(u)=\mathcal{G}_{i}^{\,\,\,k}\lim_{u\rightarrow+\infty}\Theta_{k}^{\,\,\,j}(u)=\begin{pmatrix}e^{\mathrm{i}\pi\hat{\mathcal{N}_{1}}}g_{11}&0&0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{4}}}g_{14}\\ 0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{2}}}g_{22}&e^{\mathrm{i}\pi\hat{\mathcal{N}_{3}}}g_{23}&0\\ 0&e^{\mathrm{i}\pi\hat{\mathcal{N}_{2}}}g_{32}&-e^{\mathrm{i}\pi\hat{\mathcal{N}_{3}}}g_{22}&0\\ e^{\mathrm{i}\pi\hat{\mathcal{N}_{1}}}g_{41}&0&0&-e^{\mathrm{i}\pi\hat{\mathcal{N}_{4}}}g_{11}\\ \end{pmatrix},$ (231)
and similarly at $u\rightarrow-\infty$ using (81). Now, we require the elementary property $f(u)+f^{-1}(u)=2\times\mathds{1}$, that is (80) in the symmetric sector,
$\mathcal{G}\cdot\lim_{u\rightarrow\pm\infty}\Theta\left(u\right)+\kappa\cdot\lim_{u\rightarrow\pm\infty}\Theta^{T}\left(u\right)\cdot\mathcal{G}^{T}\cdot\kappa^{-1}=2\times\mathds{1}.$ (232)
This allows certain coefficients to be expressed in terms of the charges. Now, we can impose that the $f_{i}^{j}(u)$ matrix should be built out of only four functions $\tau_{i}(u)$ (see Section 2.1.3), and at large $u$ we find
$\lim_{u\rightarrow\pm\infty}f_{i}^{\,\,\,j}(u)$ $=\left(\begin{array}[]{cccc}e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{4,\pm}+1&e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{3,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{2,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}^{2}\\ e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{4,\pm}&e^{\imath\mathcal{P}}t_{2,\pm}t_{3,\pm}+1&-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}^{2}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{2,\pm}\\ e^{\mathrm{i}\mathcal{P}}t_{3,\pm}t_{4,\pm}&e^{\mathrm{i}\mathcal{P}}t_{3,\pm}^{2}&1-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{3,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{3,\pm}\\ e^{\mathrm{i}\mathcal{P}}t_{4,\pm}^{2}&e^{\mathrm{i}\mathcal{P}}t_{3,\pm}t_{4,\pm}&-e^{\mathrm{i}\mathcal{P}}t_{2,\pm}t_{4,\pm}&1-e^{\mathrm{i}\mathcal{P}}t_{1,\pm}t_{4,\pm}\\ \end{array}\right),$ (237)
@@
and we should assume that $\tau$ has constant asymptotics in order to have a gluing matrix independent of $u$. Above, $e^{\mathrm{i}\mathcal{P}}$ is a constant appearing through (29), which in this sector satisfies $e^{2\mathrm{i}\mathcal{P}}=1$.
The crucial step is now to use the fact that
$\lim_{u\rightarrow\pm\infty}f(u)=\mathcal{G}\cdot\lim_{u\rightarrow\pm\infty}\Theta(u)$ (239)
to parametrize $\mathcal{G}$ by the charges and by the asymptotic values of $\tau_{i}(u)$. The two alternative expressions found using the relation at $u\rightarrow\pm\infty$ should coincide, since $\mathcal{G}$ is a constant. This poses some additional constraints, and they can have only two possible outcomes, either $\Delta$ or $S$ needs to be quantized to integer values. In particular, by employing the asymptotic behavior of $\Theta(u)$ along with the prototype form (230) in (237), the conditions
@@
are immediately derived. It becomes evident that the case $t_{1,\pm}=t_{4,\pm}=0$ results in the integer quantization of $\Delta$. In order to avoid this restriction, one must choose $t_{2,\pm}=t_{3,\pm}=0$. Finally, we impose the condition $\text{det}\left(\mathcal{G}\right)=1$, as discussed in Section 2.1.2. Consequently, the gluing matrix (88), as reported in the main text, directly follows. The spin will be quantized to integer values according to
$\text{e}^{\mathrm{i}\pi\hat{\mathcal{N}}_{2}}=\mp\mathrm{i},$ (241)
which is one of the constraints stemming from (239), together with
$(-1)^{M_{1}+M_{2}}=-1.$ (242)
Appendix E Some numerical results
@@
Table 1: Numerical results for the leading Regge trajectory interpolating the even spin operators in the $\Delta-S$ plane ($sl(2)$-like grading). Coupling is fixed to be $h=0.27$.
@@
$1.570786$ 0.40349193595508341712711865204630808(2)0.4034919359550834171271186520463080820.40349193595508341712711865204630808(2)0.40349193595508341712711865204630808 ( 2 )
$2.490786$ 1.15331291813913362657973321522920687(6)1.1533129181391336265797332152292068761.15331291813913362657973321522920687(6)1.15331291813913362657973321522920687 ( 6 )
$2.930786$ 1.535895832119957869450052453320828(8)1.53589583211995786945005245332082881.535895832119957869450052453320828(8)1.535895832119957869450052453320828 ( 8 )
$3.65$ 2.18041725863311622312207884034814(6)2.1804172586331162231220788403481462.18041725863311622312207884034814(6)2.18041725863311622312207884034814 ( 6 )
$4.3$ 2.77670390118381260576336103962597(5)2.7767039011838126057633610396259752.77670390118381260576336103962597(5)2.77670390118381260576336103962597 ( 5 )
$4.95$ 3.38189806050410142138805775251816(9)3.3818980605041014213880577525181693.38189806050410142138805775251816(9)3.38189806050410142138805775251816 ( 9 )
$5.7$ 4.0881362169514661658588998390465(8)4.088136216951466165858899839046584.0881362169514661658588998390465(8)4.0881362169514661658588998390465 ( 8 )
Table 2: Numerical results for the Pomeron as a function of the coupling (“bottom” grading). The lesser precision in our data for $h\geq 0.75$ is due to a choice of lower truncation parameters.
$h$ $\mathcal{S}_{0}$
$0.005$ 1.0006652065447467198180(7)1.000665206544746719818071.0006652065447467198180(7)1.0006652065447467198180 ( 7 )
$0.05$ 1.0565169798036575580020(6)1.056516979803657558002061.0565169798036575580020(6)1.0565169798036575580020 ( 6 )
$0.1$ 1.16593475741818149563694(3)1.1659347574181814956369431.16593475741818149563694(3)1.16593475741818149563694 ( 3 )
$0.15$ 1.2743232289040032420234(0)1.274323228904003242023401.2743232289040032420234(0)1.2743232289040032420234 ( 0 )
$0.4$ 1.60653570276846146957507(6)1.6065357027684614695750761.60653570276846146957507(6)1.60653570276846146957507 ( 6 )
$0.5$ 1.6736828615476622260025(2)1.673682861547662226002521.6736828615476622260025(2)1.6736828615476622260025 ( 2 )
$0.75$ $1.7736017100(7)$
$1$ $1.82742249883(1)$
$1.5$ $1.8833291288(0)$
$2$ $1.91194771316(4)$
@@
Table 3: Numerical results for the minimum of the leading odd trajectory as a function of the coupling (“bottom” grading).
$h$ $\mathcal{S}_{0}^{\text{odd}}$
$0.003$ 0.5001897150261351386032301306349(2)0.500189715026135138603230130634920.5001897150261351386032301306349(2)0.5001897150261351386032301306349 ( 2 )
$0.01$ 0.5020944885214757093188076580147(7)0.502094488521475709318807658014770.5020944885214757093188076580147(7)0.5020944885214757093188076580147 ( 7 )
$0.1$ 0.6336521170620874337755649398308(5)0.633652117062087433775564939830850.6336521170620874337755649398308(5)0.6336521170620874337755649398308 ( 5 )
$0.15$ 0.7183644121590866032768177141(3)0.718364412159086603276817714130.7183644121590866032768177141(3)0.7183644121590866032768177141 ( 3 )
$0.2$ 0.7867554536262786127683358642(8)0.786755453626278612768335864280.7867554536262786127683358642(8)0.7867554536262786127683358642 ( 8 )
$0.3$ 0.87634247722560075689570311(2)0.8763424772256007568957031120.87634247722560075689570311(2)0.87634247722560075689570311 ( 2 )
$0.5$ 0.94753956996763001(2)0.9475395699676300120.94753956996763001(2)0.94753956996763001 ( 2 )
$0.8$ $0.973009902592(3)$
$0.92$ $0.97698880504(7)$
$1$ $0.97899057845(7)$